# Tomur: Traffic-Aware Performance Prediction of On-NIC Network Functions with Multi-Resource Contention

### v3

@@ 8 Discussion
Table 9: Prediction accuracy of SLOMO and Tomur when target NF runs under memory-only contention and dynamic traffic profiles on Pensando SmartNIC.
What if the configuration of an NF changes? It is possible that an NF’s configuration is adjusted for various reasons. Such changes can cause performance characteristics to change, making the existing model inaccurate. To make Tomur “configuration-aware”, one may adopt a similar approach of traffic attributes, i.e., extracting “configuration attributes” for an NF and integrating it into the per-resource models. We leave this as future work.
9 Related Work
NF performance modeling. There have been extensive efforts on modeling NF performance. We compare Tomur with past frameworks in Table 10. Tomur is, to our knowledge, the first contention-aware framework that explicitly models multi-resource contention and traffic attributes.
@@ 9 Related Work
Clara [49]
LogNIC [32]
BubbleUp [44]
SLOMO [42]
Tomur
@@
Table 10: NF performance prediction frameworks. “Sourcecode-agnostic” means the framework does not require NF’s source code. SLOMO only considers 20% variation in flow counts.
@@
Isolation. Resource isolation techniques have been explored to provide performance guarantees of co-running applications. For example, FairNIC [31] proposes isolation solutions for SmartNIC accelerators. PARTIES [26] and ResQ [56] leverages off-the-shelf isolation techniques, e.g., Intel CAT [16] to enable QoS-aware resource partitioning. PicNIC [35] enables isolation in the network virtualization stack shared by tenant VMs. Netbricks[48] leverages its safe runtime to provide software-level isolation on servers. However, these efforts are either inapplicable to SmartNICs, or provide only partial isolation, or require substantial rewriting of NFs.
@@
SmartNIC-accelerated NFV. NFV platforms have been leveraging SmartNICs to improve energy efficiency and enable host resource-saving [41, 40, 34, 36, 51, 53]. For example, E3 [41] builds a SmartNIC-accelerated microservice execution platform with high energy efficiency. Meili[53] provides a novel framework to achieve high resource efficiency and scalable NF performance across the SmartNIC cluster. Tomur is complementary to them as it can assist operators to make better runtime decisions, thus improving resource utilization and reducing SLA violations.
10 Conclusion
@@
Prior contention-aware performance prediction frameworks fail to accurately predict the performance of on-NIC NFs due to multi-resource contention and changing traffic profiles. We systematically analyze multi-resource contention characteristics on SmartNIC as well as the impact of traffic attributes on performance of on-NIC NFs. Our insights enable the design of Tomur, an NF performance prediction framework applicable to multi-resource contention and changing traffic attributes. Tomur features per-resource contention modeling and traffic-aware argumentation to reduce multi-resource contention to single-resource contentions and integrate traffic-aware knowledge into per-resource models. Evaluations on common on-NIC NFs demonstrate that Tomur achieves accurate performance predictions, with 3.7% prediction error and 78.8% accuracy improvement on average compared to prior works, 88.5% higher resource utilization and 92.2% less SLA violations in NF placement, and 100% accurate performance bottleneck diagnosis.
Appendices
Appendix A Accuracy of Single-Resource Models
We validate the accuracy of single-resource performance model for memory subsystem contention (SLOMO) and regex accelerator contention (4.1.1), and present results in Table 11 and 12, respectively.
NF MAPE (%) $\pm$5Acc.(%) $\pm$10Acc.(%)
IPRouter 1.8 90.0 100.0
FlowStats 1.2 100.0 100.0
NAT 2.5 88.0 96.0
IPTunnel 0.6 98.0 100.0
FlowMonitor 0.9 98.0 100.0
NIDs 0.7 100.0 100.0
FlowTracker 1.6 96.0 100.0
ACL 2.2 100.0 100.0
FlowClassifier 1.1 94.0 100.0
Table 11: End-to-end prediction error of SLOMO when target NF runs under memory-only contention and fixed traffic profile. SLOMO attains high accuracy regarding memory subsystem contention at fixed traffic profile.
NF MAPE (%) $\pm$5Acc.(%) $\pm$10Acc.(%)
FlowMonitor 1.3 94.1 100.0
NIDs 1.2 100.0 100.0
Table 12: End-to-end prediction error of our regex model when target NF runs under regex-only contention and fixed traffic profile. Our regex model shows high accuracy regarding regex contention at fixed traffic profile.
Appendix B Contention in Memory Subsystem
Setup. We co-run FlowStats with mem-bench and draw NF’s throughput regarding competing working set size (WSS) at different competing cache access rate (CAR) in Figure 9. We adjust both the working set size (WSS) and cache access rate (CAR) of mem-bench to apply different types of pressure on memory subsystem.
Observation. The figure shows decreasing throughput of FlowStats with higher CAR and WSS. The performance drop is because of high cache evictions caused by the noisy neighbor’s cache access behavior. Below $\sim$6 MB competing WSS, NF performance is mainly affected by competing WSS rather than CAR, while above this threshold, CAR becomes the key actor that affects NF performance. This observation meets the 6MB size of L3 cache on a BlueField-2 SmartNIC, and also indicates the need to model the contention in SmartNIC’s memory subsystem by considering multiple fine-grained metrics of it, i.e., access rate and occupancy of last level cache.
Figure 9: The throughput of Flowstats as a function of competing working set size and cache access rate. The traffic profile has 16K flows. The packet size is 1500 B.
Appendix C Formula
Based on Equation 3 and 4, the composition function $F$ adopted in execution-pattern-based composition can represented as follows:
$F(E,T_{solo},\Delta T_{1},...,\Delta T_{r})=\begin{cases}T_{solo}-\max(\Delta T_{1},...,\Delta T_{r}),\text{if }E=pl\\ \\ \frac{1}{\displaystyle\sum_{j=1}^{r}(\frac{1}{T_{solo}-\Delta T_{j}})-\frac{r-1}{T_{solo}}},\text{if }E=rtc\end{cases}$
where $E$ represents execution pattern, $pl$ and $rtc$ stand for pipeline and run-to-completion pattern.
Appendix D Performance Counters
We list the performance counters Tomur uses to measure contention level in memory subsystem in Table 13. These performance counters are selected according to [42].
@@ Appendix D Performance Counters
Table 13: Performance counters for training the per-resource model of memory subsystem.
Appendix E Validation of Queue-Based Model
Equation 2 explicitly models NF’s performance regarding regex contention, but how it is aligned with experiment results in Figure 4 is unclear. Therefore, we show that two properties of Figure 4 — linearity and equilibrium (recall 4.1.1), can be derived and explained using Equation 2.
Linearity. In Figure 4, we run a synthetic regex-NF (denoted as $NF_{1}$) with regex-bench, each of which construct one request queue with the regex accelerator. Suppose that we control the arrival rate of regex-bench to $T_{C}$ (just as we did in Figure 4), then in each unit of time of regex accelerator, the processing of regex-bench’s request will take away $T_{C}t_{C}$ unit of time, where $t_{C}$ is the average per-request processing time of regex-bench’s requests. This means that only ($1-T_{C}t_{C}$) unit of time is left for the processing of $NF_{1}$. Therefore, the throughput of $NF_{1}$ (denoted as $T_{1}$) can be represented as:
$T_{1}$ $=(1-T_{C}t_{C})T_{1,solo}=T_{1,solo}-\frac{T_{1,solo}}{T_{C,solo}}T_{C}$ (8)
where $T_{1,solo}$ and $T_{c,solo}$ represent solo throughput of regex processing in $NF_{1}$ and regex-bench. Equation 8 is aligned with the linearity between $NF_{1}$’s throughput and competing throughput from regex-bench.
Equilibrium. Since both $NF_{1}$ and regex-bench only construct one queue with regex accelerator, we can calculate regex-bench’s equilibrium throughput as: $T_{E}=\frac{1}{\frac{1}{T_{1,solo}}+\frac{1}{T_{C,solo}}}$ based on Equation 2. By assigning $T_{E}$ to $T_{C}$ in Equation 8, we can derive that $T_{1}=(1-\frac{T_{E}}{T_{C,solo}})T_{1,solo}=T_{E}$, indicating that the regex-NF and regex-bench achieve same throughput of $T_{1}=\frac{1}{\frac{1}{T_{1,solo}}+\frac{1}{T_{C,solo}}}$. This is also aligned with the observation that an equilibrium will be achieved for NFs sharing the regex accelerator.
Appendix F Detailed Workflow of Tomur
F.1 General Workflow
To apply Tomur on an NF ($NF_{i}$), a general workflow can be split to (1) offline profiling and training, (2) online prediction.
Offline profiling and training. Tomur first collects contention level of $NF_{i}$ and synthetic competitors, and the throughput of $NF_{i}$ when co-running synthetic competitors with $NF_{i}$ under different traffic profiles. In the training phase, we train per-resource models of $NF_{i}$ using the (contention level, traffic attribute, throughput) tuple.
Online profiling. We use contention level of competitors (profiled) and traffic attributes of current traffic profile (provided) as model input to obtain . Then we compose these partial predictions based on execution pattern of $NF_{i}$ (provided) to obtain end-to-end throughput predictions.
We show details of offline profiling and training in F.2 and online prediction in F.3.
F.2 Offline Profiling and Training
Tomur collects necessary profiles of synthetic competitors and $NF_{i}$. For synthetic competitors, their contention levels on shared resources are collected. For $NF_{i}$, Tomur collects its contention level on shared resources and its throughput under different levels of single-resource contention and traffic profiles.
Contention level of synthetic competitor. We measure the contention level that synthetic competitors applies on their corresponding resource. In our current design, we consider the contention in regex accelerator and memory subsystem. Therefore, for each configuration (denoted as $c$) of synthetic competitor, we collect contention level of regex-bench ($C_{R,c}$) and mem-bench ($C_{M,c}$).
For regex-bench, we run regex-bench independently on SmartNIC. We adjust per-request processing time ($t_{R,c}$) of regex-bench by changing MTBR of input traffic profile. The number of request queues of regex-bench is fixed to 1. Then the contention level of regex-bench is represented as $C_{R,c}=(1,t_{R,c})$ for each configuration.
For mem-bench, we run mem-bench independently on SmartNIC. We adjust contention level of mem-bench by changing the memory access pattern, memory access speed and working set size of it. For each configuration, we collect runtime performance counters shown in appendix D (denoted as $pc_{1,c},...,pc_{k,c}$) of mem-bench. Therefore, the contention level of mem-bench is represented as $C_{R,c}=(pc_{1,c},...,pc_{k,c})$.
Note that the profiling of synthetic competitors is a one-time effort for all NFs, meaning that any NF can reuse the profiling data of synthetic competitors to train its models.
Contention level of $NF_{i}$. We measure the contention level that $NF_{i}$ applies on $R$ types of shared resources under different traffic profiles ($P$). A traffic profile $P$ can be denoted as a vector of traffic attributes $(p_{1},...,p_{l})$. In our current design, we measure the contention $NF_{i}$ applies on regex accelerator ($C_{i,r,P}$) and memory subsystem ($C_{i,m,P}$), respectively.
To measure $C_{i,r}$, we co-run $NF_{i}$ with regex-bench. Since the per-request processing time of regex-bench has already been profiled, we calculate the number of queues ($n_{i}$) and per-request processing time ($t_{i,P}$) of $NF_{i}$ using Equation 2. Therefore, the contention level of $NF_{i}$ on regex accelerator can be represented as $C_{i,r,P}=(n_{i},t_{i,P})$
To measure $C_{i,m}$, we run $NF_{i}$ independently on SmartNIC. We measure contention levels of $NF_{i}$ under a series of traffic profiles by collecting runtime performance counters $pc_{1,P},...,pc_{k,P}$ of $NF_{i}$. Therefore, the contention level of $NF_{i}$ on memory subsystem can be represented as $C_{i,c,P}=(pc_{1,P},...,pc_{k,P})$. We also measure the solo throughput ($T_{i,solo}$) of $NF_{i}$ by the way.
Training data of $NF_{i}$. We collect training data of $NF_{i}$ for memory subsystem model using adaptive profiling algorithm. The algorithm co-runs (1) mem-bench using different configurations ($c$) with (2) $NF_{i}$ using different traffic profile ($P$), and measure the interfered throughput ($T_{i,c,P}$) of $NF_{i}$. Note that the dimension of $P$ is pruned by adaptive profiling for each NF’s memory model.
Due to the explicit form of $NF_{i}$’s regex model, collection of a small amount of data ($<$ 10) is enough. Similarly, we co-run (1) regex-bench using different configurations ($c$) and, (2) $NF_{i}$ using different traffic profile ($P$), and measure the interfered throughput ($T_{i,c,P}$) of $NF_{i}$. Note that we only consider the change of MTBR in $P$ for regex model.
Specifically, each entry of training data for $NF_{i}$’s per-resource models can be represented as:
  1. 1.
    Memory subsystem model: $((C_{M,c},P),T_{i,c,P})$
  2. 2.
    Regex model: $((C_{R,c},P),T_{i,c,P})$
Model training. We leverage the profiled data to train per-resource models of $NF_{i}$. We denote the memory subsystem model and regex model of $NF_{i}$ as $M_{i,m}$ and $M_{i,r}$ respectively. To train $M_{i,m}$, we use $(C_{M,c},P)$ as input feature and $T_{i,c,P}$ as label. To train $M_{i,r}$, we use Equation 2 to calculate number of request queues and per-request processing time of $NF_{i}$. According to Equation 6, we use linear regression to correlate per-request processing time of $NF_{i}$ with MTBR (as in $P$).
F.3 Online Prediction
Suppose that $NF_{i}$ is co-located with a group of competitors $\mathcal{A}=\{NF_{1},...,NF_{n}\}$. To predict the throughput of $NF_{i}$, Tomur requires the following inputs:
  1. (1)
    $P$: traffic profile of $NF_{i}$ (provided).
  2. (2)
    $E$: execution pattern of $NF_{i}$ (provided).
  3. (3)
    $T_{solo,P}$: solo throughput of $NF_{i}$ using traffic profile $P$ (profiled).
  4. (4)
    $C_{j},\;\forall j\;s.t.\;NF_{j}\in\mathcal{A}$: contention level of each competitor on memory subsystem $C_{j,m}$ and $C_{j,r}$ (profiled).
We use (1) and (4) as input to $M_{i,m}$ and $M_{i,r}$ to predict throughput drop of $NF_{i}$ under traffic profile $P$ caused by contention in memory subsystem ($\Delta T_{i,m}$) and regex accelerator ($\Delta T_{i,r}$) respectively. Then we use $\Delta T_{i,m}$, $\Delta T_{i,r}$, (2) and (3) as input to our composition formula (Equation 7) to obtain end-to-end throughput prediction of $NF_{i}$.

### v5

@@ 8. Discussion
Table 9. Prediction accuracy of SLOMO and Yala when target NF runs under memory-only contention and dynamic traffic profiles on Pensando SmartNIC.
Can Yala adapt to system-wide state changes that affect the maximum performance of NFs? Yala predicts the maximum throughput of co-located NFs to ensure they conform to SLA under multi-tenancy. However, SmartNICs may experience system-wide state changes due to various reasons, altering the maximum throughput of NFs. A typical example is dynamic voltage and frequency scaling (DVFS) (dvfsarm, ; dvfs10, ), which may affect the NF’s maximum performance due to CPU frequency down-scale or up-scale. Different DVFS policies can be enabled by setting the frequency scaling governor of CPUs on servers, which is however not supported on common SoC SmartNICs currently, e.g. the SmartNICs we use in our experiments (Nvidia Bluefield-2 and AMD Pensando) (bfscripts, ). If we are to extend Yala to predict maximum performance of on-NIC NFs on future generations of SmartNICs that support DVFS and other similar system-wide state changes, related variables, e.g. OS policy and thermal state for DVFS, should be integrated in our models to ensure high accuracy under different system states, e.g. power state in DVFS.
9. Related Work
NF performance modeling. There have been extensive efforts on modeling NF performance. We compare Yala with past frameworks in Table 10. Yala is, to our knowledge, the first contention-aware framework that explicitly models multi-resource contention and traffic attributes.
@@ 9. Related Work
Clara (clara21, )
LogNIC (lognic23, )
BubbleUp (bubbleup11, )
SLOMO (slomo20, )
Yala
@@
Table 10. NF performance prediction frameworks. “Sourcecode-agnostic” means the framework does not require NF’s source code. SLOMO only considers 20% variation in flow counts so it is considered “half” traffic-aware.
@@
Isolation. Resource isolation techniques have been explored to provide performance guarantees of co-running applications (fairnic20, ; parties19, ; resq18, ; picnic20, ; netbricks, ; osmosis24, ). For example, FairNIC (fairnic20, ) proposes isolation solutions for SmartNIC accelerators. PARTIES (parties19, ) and ResQ (resq18, ) leverages off-the-shelf isolation techniques, e.g., Intel CAT (intelcat, ) to enable QoS-aware resource partitioning. However, these efforts are either inapplicable to SmartNICs, or provide only partial isolation, or require substantial rewriting of NFs.
@@
SmartNIC-accelerated NFV. NFV platforms have been leveraging SmartNICs to improve energy efficiency and enable host resource-saving (e319, ; ipipe19, ; exoplane23, ; uno17, ; db23, ). For example, E3 (e319, ) builds a SmartNIC-accelerated microservice execution platform with high energy efficiency. Yala is complementary to them as it can assist operators to make better runtime decisions, thus improving resource utilization and reducing SLA violations.
10. Conclusion
@@
Prior contention-aware performance prediction frameworks fail to accurately predict the performance of on-NIC NFs due to multi-resource contention and changing traffic profiles. We systematically analyze multi-resource contention characteristics on SmartNIC as well as the impact of traffic attributes on performance of on-NIC NFs. Our insights enable the design of Yala, a multi-resource contention-aware and traffic-aware performance prediction framework for on-NIC NFs. Yala achieves accurate performance predictions, with 3.7% prediction error and 78.8% accuracy improvement on average compared to prior works, and enables new usecases.
Appendix A Artifact Appendix
A.1. Abstract
This artifact contains source code and related tools for Yala, a multi-resource contention- and traffic-aware performance prediction framework for on-NIC NFs. Yala is publicly available on Github (yalagithub, ). Specifically, we provide source code of model training and prediction. Example profiles of an NF (FlowMonitor) can be used to train its models and produce predictions on its throughput for demonstration. In addition, we open-source the benchmark NFs (mem-bench, regex-bench and compression-bench), real NFs, NF framework (a modified version of Click), rulesets and related tools mentioned in our paper.
A.2. Artifact check-list (meta-information)
  • Compilation: We provide python scripts for training and prediction, which do not require a compiler. Compiling Click and NFs requires gcc (click2000, ). Compiling rulesets requires RXP compiler. Makefiles are provided for compilation.
  • Binary: We provide python scripts for training and prediction instead of binaries. Some binaries of related tools are included, e.g., compiled ruleset.
  • Model: Gradient boosting regression and linear regression from sklearn (sklearn, ).
  • Hardware: We use NVIDIA BlueField-2 MBF2H332A-AENOT SmartNIC (bf21, ). The performance counters in Table 11 can be accessed on this SmartNIC.
  • Publicly available?: Yes, on https://github.com/NetX-lab/Yala
  • Code licenses?: BSD 3-Clause ”New” or ”Revised” License.
@@ A.2. Artifact check-list (meta-information)
Table 11. Performance counters for training the per-resource model of memory subsystem.
A.3. Description
A.3.1. How to access
Yala is publicly available on Github: https://github.com/NetX-lab/Yala.
A.3.2. Hardware dependencies
We profile NFs on NVIDIA BlueField-2 MBF2H332A-AENOT SmartNIC. The traffic generator uses another ConnectX-6 100GbE NIC. The training and prediction do not have specific hardware requirements.
A.3.3. Software dependencies
Dependencies of the software are listed as below:
  • Training and prediction
    • Python: 3.8
    • scikit-learn (sklearn, ): 0.24.2
    • numpy: 1.19.5
    • pandas: 1.1.5
    • tabulate: 0.9.0
  • Traffic generator
  • NF frameworks
A.4. Installation and Testing
For installation of hardware and software dependencies, please follow our official guide on Github. The training and prediction of Yala can be tested upon satisfying software dependencies.
We provide an example of training and using the model to predict throughput for FlowMonitor. To train Yala using example training set, please go to /model directory and run python3 train.py command. This generates a models.pkl file containing linear model, memory-only model (SLOMO), regex-only model and Yala’s model for FlowMonitor as an example. Then to evaluate the model accuracy on the example test set, run python3 predict.py command. This reports MAPE, ±$\pm 5\%$ Acc. and ±$\pm 10\%$ Acc of the four models mentioned.
A.5. Other Notes
Detailed instructions on using open-sourced tools, e.g., benchmark NFs, can be found in ”additional tips” on our Github.
A.6. Methodology
Submission, reviewing and badging methodology: